OpenStudy (anonymous):

factorize completely 81_m^2

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\huge 81m^2$$

OpenStudy (skullpatrol):

$$\huge (ab)^m=a^m*b^m$$

OpenStudy (skullpatrol):

$$\huge 9*9=81$$

OpenStudy (jdoe0001):

hmmm \(\huge 81-m^2\quad ?\)